A sphere is charged to 20 µC. Find the potential at the following points. A, B and C (A
point is at the center of the sphere, B is at the surface and C is 4cm from the surface.)
1
Expert's answer
2020-08-12T16:19:14-0400



1) Let's choose the spherical shell inside the charged shere (drawn with thin line). Then, according to the Gauss's theorem, the flux of electrical field through this shell will be:


"\Phi =ES = 4\pi k q_{inside}"

where "S" is the area of the shell surface, "k = 9\times 10^9 N\cdot m^2/C^2" is the constant, and "q_{inside}" is the charge inside the shell. Since the sphere is empty ( "q_{inside} = 0" ), then "E = 0" in any point inside the shere. Particulary, "E_A = 0."

As far, as the potential is connected with the electric field in the following way:


"E = -\dfrac{d\varphi}{dr}"

the potintial in the point "A" will be constant "\varphi_A = const".


2) Let's choose the spherical shell outside the charged shere (drawn with thin line) that passes through the point "C".

Then, according to the Gauss's theorem, the flux of electrical field through this shell will be:

"\Phi =E_CS = 4\pi k q"

where "S = 4\pi r^2" is the area of the shell surface and "q = 20\times 10^{-6}C" is the charge enclosed by this surface (the charge of the thick sphere). Hence, obtain:


"E_C\cdot 4\pi r^2 = 4\pi kq\\
E_C = k\dfrac{q}{r^2}"

The potential will be:


"\varphi_C = -\int E_C dr = -\int k\dfrac{q}{r^2}dr = k\dfrac{q}{r} + C"

we can choose the contstan "C" to be zero if we set the potential be zero at large distance from the sphere. Thus:


"\varphi_C = k\dfrac{q}{r} = k\dfrac{q}{R+4cm}"

The distance "r" can be considered as "r = R+4cm", where "R" is the radius of the charged sphere.

3) Then, the potential on the sphere surface is:


"\varphi_B = k\dfrac{q}{R}"

where "R" is the radius of sphere.

Since the potential should be continuous, the potential inside the sphere should be equal to the potential on the surface. Thus:

"\varphi_A = const = \varphi_B = k\dfrac{q}{R}"

4) Let's summarize the results:


"\varphi_A = k\dfrac{q}{R}\\
\space\\
\varphi_B=k\dfrac{q}{R}\\
\space\\
\varphi_C=k\dfrac{q}{R+4cm}\\"

As far as the radius "R" of the charged sphere is unknown, we can not give the final numerical answer. But let's come as close as we can:

"\varphi_A = 9\times 10^9\cdot \dfrac{20\times 10^{-6}}{R} = 1.8\times10^5 \cdot \dfrac{1}{R}\space [V]\\
\space\\
\varphi_B=1.8\times10^5 \cdot \dfrac{1}{R}\space [V]\\
\space\\
\varphi_C=1.8\times10^5 \cdot \dfrac{1}{R+0.04m}\space [V]\\"

Substituting the radius "R" of the charged sphere in meters, obtain the final answer.


Answer.

"\varphi_A = 1.8\times10^5 \cdot \dfrac{1}{R}\space [V]\\
\space\\
\varphi_B=1.8\times10^5 \cdot \dfrac{1}{R}\space [V]\\
\space\\
\varphi_C=1.8\times10^5 \cdot \dfrac{1}{R+0.04m}\space [V]\\"